State a) work-energy theorem in rotational motion and b) momentum im
Answer #1

a)Work energy theorem for rotational motion states that the change in kinetic energy of an object is equal to the net work done by all the torques on the object executing rotational motion

W(torque) = 1/2 I wf^2 - 1/2 I wi^2

b)The impulse moment theorem states that the impulse is equal to the change in momentum of the object, when it undergoes a change in momentum and it has a force applied for a certain amount of time.

J = m (v2 - v1) ; J = F t